OpenStudy (anonymous):

Thre people in an election split the vote in the ratio of 5 to 4 to 2. The number of votes cast was 55,110. How many votes did each person receive.

OpenStudy (anonymous):

Let the number of people be 5x,4x and 2x then 5x+4x+2x=55110

OpenStudy (anonymous):

at the end however u have to multiply x by 2, 4 and 5 for each person respectively to get their full amount of votes

OpenStudy (anonymous):

Yeap

OpenStudy (anonymous):

Is 25050,20040,10020 Right?

OpenStudy (anonymous):

yeap

OpenStudy (anonymous):

Thanks both of u guys

OpenStudy (anonymous):

5x+4x+2x = 55110 11x = 55110 x = 55110/11 x = 5010 5x = 5*5010 = 25050 4x = 4*5010 = 20040 2x = 2*5010 = 10020 25050:20040:10020

OpenStudy (anonymous):

np